OpenStudy (anonymous):

Jose has $2.50 in nickels and dimes in his jar. If he has 11 more nickels than dimes, how many nickels does Jose have in his jar?

zepdrix (zepdrix):

If we let \(\large d\) be dimes and \(\large n\) be nickles The value of the total number of dimes would be \(\large .1d\) (That's 10 cents times each dime we have). If we do the same with the nickles \(\large .05n\) (5 cents times the number of \(\large n\) nickles) If all of his nickles and dimes add up to $2.50, we can write it like this. \(\large .1d+.05n=2.50\) Some number of nickles and dimes at their given values add up to 2.50.

zepdrix (zepdrix):

They gave us a little more information, `he has 11 more nickels than dimes` Which we can write this way, \(\large d+11=n\) How ever many dimes he might have, his nickles will add up to 11 more than that.

OpenStudy (anonymous):

Ok I underrstand up to here

zepdrix (zepdrix):

Let's rewrite this equation in terms of nickles. So we'll subtract 11 from each side.\[\large d+11=n \qquad \rightarrow \qquad \color{green}{d=n-11}\] We want to use both of these equations in order to solve for \(\large n\), the number of nickles. So we'll use the green equation above, and plug it into the other equation we figured out.\[\large .1\color{green}{d}+.05n=2.50 \qquad \rightarrow \qquad .1\color{green}{(n-11)}+.05n=2.50\]

OpenStudy (anonymous):

so do i solve this and It will show the answer to the number of nickels

OpenStudy (anonymous):

the number of dimes is 13 correct?

zepdrix (zepdrix):

Yes, with the way we setup the problem, we let \(\large n\) represent the number of nickles in the jar. So if we can solve for \(\large n\), the problem is solved :)

zepdrix (zepdrix):

dimes? ummmm yah that sounds correct :)

zepdrix (zepdrix):

If the number of dimes is 13, we can plug that into this equation. \(\large d=n-11 \qquad \rightarrow \qquad 13=n-11\) to solve for n.

OpenStudy (anonymous):

24!!

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

Thanks